OpenStudy (anonymous):

would \[LT \left[ tu(t-30) \right]=\frac{ e ^{-30s} }{ s ^{2} }\] i'm not sure if applying the formula below would be right. \[LT \left[ u _{c}(t)f(t-c) \right]=e ^{-cs}F(s)\]

OpenStudy (anonymous):

we know that \[LT \left[ u _{c}(t)=u(t-c) \right]=\frac{ e ^{-cs} }{ s }\]

OpenStudy (anonymous):

just from laplace tables

OpenStudy (anonymous):

ideas?

OpenStudy (anonymous):

Have you tried going through the integration?

OpenStudy (anonymous):

im unsure how to do that when we are considering u(t-30)

OpenStudy (anonymous):

Do you know how to work with the dirac delta function?

OpenStudy (anonymous):

not really, we just have a table of lapalce transforms used to help us in process control. my laplace transform is limited considering we weren't taught this in engineering maths yet it has come up in a third year topic which applies laplace transforms

OpenStudy (anonymous):

like, i start with the integral definition, \[LT(f(t)=\int\limits_{0}^{\infty}e ^{-st}tu(t-30)dt\]

OpenStudy (anonymous):

and try integration by parts...but i'm stuck with the t*u(t-30) bit, with applying the product rule on it

OpenStudy (anonymous):

Ok, so if you are using integration by parts, then you are considering this integral of the from\[\int\limits_{}^{}u dv = uv - \int\limits_{}^{}v du\]. What are you using as your u and your dv?

OpenStudy (dan815):

do a change of variables

OpenStudy (anonymous):

dv is the exponential and u is the t*u(t-30)

OpenStudy (anonymous):

ohh yea neat

OpenStudy (anonymous):

Thanks @dan815

OpenStudy (anonymous):

yeah nah that drawing has confused me lel

OpenStudy (dan815):

lol

OpenStudy (dan815):

i realized using s for the different variable was bad idea lol

OpenStudy (anonymous):

haha no joke i was gonna say

OpenStudy (dan815):

now u became the variable for laplace

OpenStudy (dan815):

lemme think so i wonder if we gotta add some extra stuff in for this shift

OpenStudy (anonymous):

could we just expand those brackets

OpenStudy (anonymous):

so instead of s lets use 'a' as our variable

OpenStudy (anonymous):

(a+30)u(a)= au(a)+30u(a)

OpenStudy (dan815):

its best to jst graph this stuff so we know for sure we dont have anything funky going on

OpenStudy (anonymous):

and laplacing these would be the same...

OpenStudy (anonymous):

where a=t-30

OpenStudy (dan815):

ohh something silly here oops

OpenStudy (dan815):

ont need to even worry about all this apprently

OpenStudy (anonymous):

lel

OpenStudy (dan815):

this is U_c(t) by definition

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but we can't use the formula thats on the top page cause it involves f(t-c) so here c=0 for the function f, but 30 for variable 'u'

OpenStudy (dan815):